Box going up elevator uniformly at 5m/'s falls off and hits bottom in 1.6secs

A) what time to reach max height?
B) height at which it fell?
C) height at 0.2 secs after falling
1
Expert's answer
2019-04-10T09:42:23-0400

b) First we find the height from which the box fell. Use the formula

"h\left( t \right)=\frac{g{{t}^{2}}}{2}"

where g=9.8 m/s2 is the acceleration of gravity, t=1.6 secs.Substituting these values, we get


"h=\frac{9.8\cdot {{1.6}^{2}}}{2}=12.5\,\text{m}"

So the height at which the box fell is h=12.5 meters.

a) Now we find the time for which the box reached its maximum height. As the box going up elevator uniformly at v=5 m/s the height to which it will rise during time t is


"h=v\cdot t"

Substitute h=12.5 m and v=5 m/s


"12.5=5\cdot t"

Solving this for t we get the time for which the box reached its maximum height


"t=\frac{12.5}{5}=2.5\,s"

c) To find height at 0.2 secs after falling first we find the distance that the box will cover in 0.2 seconds


"s\left( 0.2 \right)=\frac{g{{t}^{2}}}{2}=\frac{9.8\cdot {{0.2}^{2}}}{2}=0.2\,\text{m}"

and then subtract this distance from the height of the fall. We get


"h\left( 0.2 \right)=12.5-0.2=12.3\,\text{m}"

So height at 0.2 secs after falling is 12.3 m